OpenStudy (anonymous):

1, 1, 2, 3, 5, 8 ___.

OpenStudy (anonymous):

that looks like the Fibonacci sequence so it would be 13. 1+1=2 1+2=3 2+3=5 3+5=8 etc

OpenStudy (anonymous):

ohhh great..! excellent.